OpenStudy (anonymous):

how to do the derivative of sqrt(t)*(1-t^2)

OpenStudy (anonymous):

Use product rule.

OpenStudy (anonymous):

\[ (\sqrt{t}(1-t^2))'=(\sqrt{t})'(1-t^2)+(\sqrt{t})(1-t^2)' \]

OpenStudy (anonymous):

after multiplying from both sides. what do I do about the powers.

OpenStudy (anonymous):

Differentiate the parts with in \((\ldots )'\)

OpenStudy (anonymous):

The parts without \('\) aren't differentiated.

OpenStudy (anonymous):

what I have right now is (1-5t^(3/2)/2t^(1/2). does need it to change further?

OpenStudy (anonymous):

I'm not really sure how you got that.

OpenStudy (anonymous):

What I get is \[ \frac{1-t^2}{2\sqrt{t}}-2t\sqrt{t} \]

OpenStudy (anonymous):

Though I didn't simplify.

OpenStudy (anonymous):

Simplifying a step gives\[ \frac{1-t^2-4t^2}{2\sqrt{t}}=\frac{1-5t^2}{2\sqrt{t}} \]